OpenStudy (anonymous):

Let A = 40. If B is complement of A, and C is a supplmenet of B, what is B+C? A. 60 B. 150 C. 180 D. 210

OpenStudy (ipwnbunnies):

Do you know what complementary and supplementary angles are?

OpenStudy (anonymous):

Oh wait.. Complementary angles are angles that add up to 90 degrees. Supplementary angles are angles that add up to 180 degrees.

OpenStudy (ipwnbunnies):

Exactly.

OpenStudy (ipwnbunnies):

So, let's do the first part. It says B is a complement of A. Can you find B?

OpenStudy (anonymous):

A is 40, so B is = 40?

OpenStudy (anonymous):

or 50?

OpenStudy (anonymous):

They add to 90

OpenStudy (ipwnbunnies):

No, since A and B are complements, remember, they have to add up to 90 degrees. A + B = 90 40 + B = 90 So, yes, B has to be 50 degrees.

OpenStudy (anonymous):

So, now how do we find C?

OpenStudy (ipwnbunnies):

It says C is the supplement of B. So, B + C = 180 degrees

OpenStudy (anonymous):

C = 150

OpenStudy (ipwnbunnies):

Err, not quite. Check out math again. o-o We found B to be 50 degrees

OpenStudy (anonymous):

130

OpenStudy (anonymous):

Oh, I'm slow. I already solved the question. It asked what they were equal to which is 180.. LOL

OpenStudy (ipwnbunnies):

Mhm. c: Ohno, I just realized the end of this question. xD But it's practice. It just ask for B+C, but we could've known it was 180 from the start.

OpenStudy (ipwnbunnies):

I KNOW RIGHT LOLOLOL

OpenStudy (anonymous):

Thanks man